OpenStudy (anonymous):

How many units apart are the points (4, -3) and (-4, 3)?

OpenStudy (jdoe0001):

\(\bf \text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (anonymous):

I got 5.29 is that correct? @jdoe0001

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\\ \textit{(4, -3), (-4, 3)}\\ d = \sqrt{((-4)-(4))^2 + ((3)-(-3))^2} \implies \sqrt{(-8)^2+(6)^2}\)

OpenStudy (anonymous):

(-8)2= -64 (6)2=36

OpenStudy (jdoe0001):

\(\bf (-8)^2 \implies -8 \times -8\)

OpenStudy (anonymous):

64

OpenStudy (anonymous):

square root of 96 is 9.79

OpenStudy (jdoe0001):

\(\bf d = \sqrt{((-4)-(4))^2 + ((3)-(-3))^2} \implies \sqrt{(-8)^2+(6)^2}\\ \implies \sqrt{64+36}\)

OpenStudy (anonymous):

64+36=96

OpenStudy (anonymous):

I mean 100

OpenStudy (anonymous):

oh its 10

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

I got it thanks

OpenStudy (jdoe0001):

yw